problem 2-3 Engineering Study Advancement Committee planned a summer training program with following data for School administrative approval: 1. What is the program cost break-even point (how many students)? 2. If 40 students registered for this training, school will open another classroom. what is the marginal cost? 3. How many students required for this program profitable? food and cost formaterial perstudent revenue per student siz4classroostudent 225 225 classroomA classroom B student registration 0 to 40 35
Answer #1
  1. Break Even Point = Fixed Cost / Sales Price Per Unit – Variable Cost per Unit

OR Fixed Cost / Contribution Margin Per Unit

In this above case,

Fixed Cost = Cost for Classroom i.e. 225

Sales Price per Unit (Revenue per student) = 35

Variable Cost per unit (Food and material per student) = 20

                Break-Even Point = 225/ (35-20) = 225/15=15

  1. Marginal Cost: If >40 student registered for the training, school will open another classroom, Marginal cost is the additional cost incurred for the production of an additional unit of output. In our case

Marginal cost = Change in total cost of program/ change in registered student by unit 1

Change in total cost of program when 41st student registers for the training = Cost of Additional classroom (225) + cost of Food and Material for additional student (20)

Marginal Cost = 225+20/1 = 240

  1. Students Required for the program to be profitable when students are >40

Contribution margin per student = Revenue per student – Variable cost per student (food and material cost)

= 35-20 = 15

No. of students required for making the program profitable = Additional fixed cost /Contribution margin = 225/15 = 15

Total students required to make the program profitable when 2 classrooms are used = 15+15 = 30